# Computing Anharmonic Infrared Spectra of Polycyclic Aromatic Hydrocarbons Using Machine-Learning Molecular Dynamics

### v2

@@ Abstract
Polycyclic aromatic hydrocarbons (PAHs) are key contributors to interstellar aromatic infrared (IR) bands. However, current spectral databases for IR emission analysis are limited by the omission of vibrational anharmonicity and temperature effects, primarily because of the high computational cost of conventional quantum chemical calculations (QCCs). In this work, we present a machine learning-based molecular dynamics (MLMD) approach that efficiently computes anharmonic IR spectra while incorporating temperature effects. MLMD achieves predictive accuracy comparable to that of QCCs but with significantly reduced computational cost, scaling linearly with the number of atoms in the system. We applied MLMD to calculate the anharmonic spectra of 1 704 PAHs in the NASA Ames PAH IR Spectroscopic Database with up to 216 carbon atoms, demonstrating its capability for high-throughput spectral calculations of large molecular systems. Our results highlight MLMD’s potential to enable the development of extensive molecular spectral datasets, enhancing data-driven analyses of astronomical IR spectra, particularly in anticipation of upcoming data from the James Webb Space Telescope.
astronomical data bases: miscellaneous; software: data analysis; ISM: molecules; infrared: ISM
I Introduction
@@
In the mid-1970s, distinct infrared (IR) signals were identified in the universe, notably at wavelengths of 3.3, 6.2, 7.7, 8.6, 11.2, and 12.7 $\upmu$m. These signals were later linked to the vibrations of C-H and C-C bonds in aromatic molecules, known as aromatic IR bands (AIB) (Leger & Puget, 1984; Allamandola et al., 1985). Such emissions are primarily attributed to polycyclic aromatic hydrocarbons (PAHs) after the absorption of ultraviolet (UV) photons. PAHs are recognized as key contributors to the evolution of the interstellar medium, influencing processes such as gas heating, ionization balance, and star formation (Tielens, 2008). Determining the population, composition, size distribution, charge state, and chemical structures of interstellar PAHs provides valuable insight into the physical conditions of their host galaxies. However, this remains a significant challenge because of the structural diversity of PAHs. Variations in carbon-ring arrangements, side groups, substitutions, charge states, and isotopologues lead to distinct IR emission characteristics, each governed by specific vibration modes (Peeters et al., 2021).
@@
In order to tackle these challenges, quantum chemical calculations (QCCs) are now an indispensable aid in AIB analysis. Using QCCs, publicly available spectral databases have been established for AIB analysis, such as the NASA Ames PAH IR Spectroscopic Database (PAHdb) (Mattioda et al., 2020; Bauschlicher et al., 2018; Boersma et al., 2014). The availability of these databases enables extensive data-driven analysis of AIBs, providing a more robust interpretation of their features (Boersma et al., 2013, 2015; Maragkoudakis et al., 2020; Sadjadi et al., 2015; Cami, 2011; Shannon & Boersma, 2019; Ricca et al., 2021). Furthermore, these databases have facilitated the analysis of astrochemical components using machine learning (ML) models, demonstrating significant computational efficiency (Kovács et al., 2020) and helping to uncover relationships between molecular structures and spectra (Meng et al., 2021, 2023). Despite these advancements, current theoretical spectral databases face three primary limitations:
@@
  • Harmonic oscillator approximations: Most spectra neglect anharmonic vibrations, leading to significant discrepancies with experimental observations (Mackie et al., 2015, 2018a; Maltseva et al., 2016, 2018; Lemmens et al., 2019).
  • Neglect of temperature effect: AIB features are typically attributed to IR emission from far-UV-pumped PAHs (Leger & Puget, 1984), where an absorbed UV photon can elevate the effective molecular temperature to several hundred or even thousands of kelvin, an effect that is not taken into account in most theoretical spectral data.
  • Predominantly featuring small molecules: For example, a significant portion of PAHs in PAHdb consists of fewer than 35 carbon atoms, while larger PAHs ($\sim$50 carbon atoms) are more likely sources of AIBs (Sellgren, 1984; Allamandola et al., 1989; Bauschlicher et al., 2009; Maragkoudakis et al., 2020)
@@
These challenges arise mainly from the high computational cost of QCCs. In particular, the computation time for anharmonic molecular spectra using second-order vibrational perturbation theory (VPT2) is prohibitively long, typically only accommodating molecules with $N_{\text{C}}<25$, where $N_{\text{C}}$ represents the number of carbon atoms in the system (Mackie et al., 2016; Esposito et al., 2024a). The cost of incorporating temperature effects into these calculations using the Wang-Landau random walk algorithm is even more prohibitive, being orders of magnitude higher due to the need for extensive sampling (Chen, 2018; Chen et al., 2019). Facing the vast diversity of PAH species, these limitations significantly hinder the accurate interpretation of observational data, posing substantial challenges in the era of advanced observational facilities such as the James Webb Space Telescope (JWST) (Boersma et al., 2023; Ricca et al., 2024).
@@
As a result, there is an urgent need for a cost-effective and accurate method that incorporates both anharmonic and temperature effects in the construction of spectral databases to improve our understanding of the origin of AIBs. In this work, we propose a ML-based molecular dynamics (MLMD) approach for computing PAH IR spectra. This method replaces computationally expensive electronic structure calculations with far more efficient ML-based ones. More importantly, it explicitly accounts for anharmonic vibrations and temperature effects. We show that the MLMD approach can reproduce experimental spectra with accuracy comparable to state-of-the-art VPT2 QCCs, while significantly reducing computational time, achieving a scaling approximately linear in $N_{\text{C}}$.
II Methods
@@
Figure 1: Schematic representation of the workflow for computing anharmonic IR spectra.
@@
II.1 Training Data
The chemical structures of 687 neutral PAH molecules were extracted from the theoretical dataset of PAHdb (v3.2, Mattioda et al. (2020)) to construct the training datasets. The selected PAHs encompass a diverse array of chemical structures that are of astronomical significance. Classical MD simulations were performed using a custom code to generate random configurations of the atoms in each molecule at 300 K, using the adaptive interatomic reactive empirical bond order potential (Stuart et al., 2000). In these simulations, the molecules reached thermal equilibrium in 200 ps using a canonical Nose-Hoover thermostat, with a time step of 0.5 fs. Ten extended structures were extracted from the atomic trajectories of each molecule. As these geometries do not necessarily fall near equilibrium configurations, random Gaussian perturbations were applied to the ground-state atomic coordinates to create 15 more configurations for each molecule to improve the generalizability of the model. Finally, in total, 17 175 configurations (also called molecular conformations) were generated.
Subsequently, single-point electronic structure QCCs were performed for each of these 17 175 configurations to calculate the potential energy, the forces in the nuclei, and the dipole $\mathbf{p}$, within the framework of density functional theory (DFT) at the B3LYP/4-31G level, as implemented in the Gaussian 16 software package (Frisch et al., 2016). This level of theory offers a favorable balance between accuracy and computational efficiency for large PAHs (Stephens et al., 1994), given its extensive application in PAH IR spectrum studies (Bauschlicher & Bakes, 2000; Ricca et al., 2012). We note that the basis set does not include polarization functions, which may introduce more uncertainties than larger basis functions (e.g 6-31g*) for PAHs containing nitrogen (Ricca et al., 2021, 2024). However, this choice was made to ensure computational efficiency, which is crucial for large-scale sampling. Through these calculations, we generated two datasets for training the NNFF and EPNN models, respectively. The first dataset comprises 17 175 energy and 2 274 825 force data points, while the second dataset includes 17 175 molecular dipole moments.
II.2 NNFF
@@
II.3 EPNN
@@
II.4 Anharmonic IR spectrum
To compute the anharmonic IR spectrum of a PAH, MD simulations were performed to simulate its finite-temperature dynamics based on atomic forces predicted by the trained NNFF. We used the MD implementation in the Atomic Simulation Environment (ASE) (Larsen et al., 2017). Initially, the molecule reaches thermal equilibrium in 200 ps in a canonical ensemble (NVT) at a target temperature controlled by the Nose-Hoover thermostat, with a timestep of 0.5 fs. The system was then further simulated in a microcanonical ensemble (NVE) for 200 ps, recording atomic configurations at every 1.0 fs during vibration.
These configurations serve as input to the EPNN, which is used to compute $\mathbf{p}(t)$ at each time step of the vibration. The dipole time autocorrelation function $\langle\mathbf{p}(0)\cdot\mathbf{p}(t)\rangle$, characterizing the time evolution of $\mathbf{p}$, is then subjected to a Fourier transform to obtain the IR intensity(berens1981):
@@ II.4 Anharmonic IR spectrum
where $\omega$ denotes the angular frequency, $k_{\text{B}}$ represents the Boltzmann constant, $\hbar$ is the reduced Planck constant, $c$ is the speed of light, $T$ indicates the temperature, and $t$ is time.
II.5 Reference methods for comparison
@@
The 49 PAHs for which the experimental spectra are available in PAHdb (v3.2), ranging from 10 to 50 carbon atoms, exhibit a variety of chemical structures, as detailed in Appendix I. The experimental IR spectra were obtained using matrix isolation techniques at low temperatures (Hudgins & Allamandola, 1995; Hudgins & Sandford, 1998a, b, c; Hudgins & Allamandola, 1999; Mattioda et al., 2003, 2005, 2014, 2017). To facilitate comparison with experimental data, we employed theoretical spectra computed using MLMD at an effective temperature of 50 K, which accounts for molecular vibrations under experimental conditions (Esposito et al., 2024b). In the comparison between computed and experimental spectra, discrete infrared spectral lines were Lorentzian broadened with a full width at half maximum of $18\text{cm}^{-1}$.
III Results and discussion
@@
Figure 3: Comparison of IR spectra for five PAHs computed using the MLMD at 50 K(left, red), DFT harmonic (middle, green), and DFT anharmonic (right, blue) methods, alongside the experimentally measured spectra (black). The intensities are normalized to the maximum values across the four spectra for ease of comparison. The RMSE is computed from the original intensity values, with units of $10^{5}\text{cm}^{2}/\text{mol}$.
@@
Figure 4 presents the distribution of errors in predicting the IR spectra of the 49 PAHs relative to the experimental data, for each of the three computational methods. The results clearly demonstrate that the harmonic spectra exhibit significantly poorer agreement with the experimental data compared to the anharmonic spectra, consistent with the findings of previous studies (Mackie et al., 2015; Maltseva et al., 2016; Lemmens et al., 2019). The precision of the MLMD and DFT anharmonic methods is comparable, with mean RMSE values of 0.363 and 0.360 ×$\times 10^{5}\text{cm}^{2}/\text{mol}$, respectively. For PAHs with fewer than 16 carbon atoms, the DFT anharmonic method generally outperforms MLMD. However, as the size of the PAHs increases, MLMD tends to provide more accurate predictions than DFT anharmonic.
Figure 4: Distribution of RMSE values for spectra predicted by the three computational methods, compared to the experimental spectra of the 49 PAHs. $N_{\text{C}}$ represents the number of carbon atoms in each molecule. Note that the anharmonic DFT calculation is not performed for large PAHs (Section II.5).
@@
To demonstrate the precision, we compared theoretical anharmonic spectra with experimental IR spectra of two PAHs, pyrene and benzo[$k$]fluoranthene, measured at different temperatures. The higher temperature gas-phase experimental spectra have been taken from the NIST Chemistry WebBook database (Linstrom & Mallard, 2001), while the lower temperature spectra come from matrix isolation experiments of Hudgins & Sandford (1998a); Hudgins & Allamandola (1999). Figure 6 shows the NIST spectra (gray curves), referred to as $T_{\text{high}}$, and the low-temperature spectra (black curves), referred to as $T_{\text{low}}$. Comparison of the two sets of experimental spectra measured at different temperatures reveals that thermal effects lead to notable changes in spectral characteristics, such as band broadening and peak shifts (indicated by dashed lines), which aligns with previous observations of Chen (2018).
@@
The two pairs of experimental spectra shown in the top panels of Figure 6 a and b are compared with three theoretical spectra in the bottom panels, calculated using MLMD at 300 (red) and 50 K (blue), as well as by the anharmonic DFT method (green). Compared with the 50 K MLMD results, the temperature-induced band broadening and peak shift behaviors are better captured by the MLMD results at 300 K, with a low RMSE value of 0.060 for the two molecules. In contrast, the agreement with the room temperature experimental spectra is poorer for both the 50 K MLMD and anharmonic DFT results, with higher RMSE values of 0.098 and 0.072 for pyrene, and 0.069 and 0.062 for benzo[$k$]fluoranthene, respectively.
@@
Despite the growing need to incorporate anharmonicity and temperature effects into astronomical spectroscopic analyses, driven by high-resolution infrared observations in the JWST era (Peeters et al., 2021), only a limited number of theoretical anharmonic spectra for PAHs, likely fewer than 100, are available in the literature. Typically, these studies focus on just 2-5 PAH species, most of which contain fewer than 20 carbon atoms (Esposito et al., 2024c, d; Mackie et al., 2022, 2021). This narrow scope makes it difficult to draw general conclusions, especially given the vast diversity of PAH chemical structures. Such limitations hinder the accurate interpretation of AIBs and are mainly due to the high computational cost of traditional QCCs. To address this gap, we have computed the anharmonic spectra at 50, 300, and 600 K for 1 704 molecules from version 3.20 of PAHdb, which includes neutral PAHs as large as 216 carbon atoms. This dataset is open to the public as described in the Section of Data Availability.
@@
This dataset is expected to be a valuable resource for the spectral decomposition of AIBs and for data-driven studies investigating the precise structure-spectral relationships of PAHs. In addition, it will aid in the exploration of temperature effects. For instance, Figure 7 displays the anharmonic IR spectra of a simple mixture of the 1 704 neutral PAHs at 50 K (blue), 300 K (green), and 600 K (red). These spectra represent the sum of individual PAH spectra and are compared to the ground-state harmonic spectra from PAHdb (gray). Notably, the IR signals are significantly enhanced at high temperatures, exhibiting stronger peaks and an elevated plateau. A redshift in peak positions is also observed at high temperatures as a general trend. Furthermore, some emission bands that are weak at lower temperatures become significantly stronger at 300 and 600 K. For example, a distinct band emerges in the 3.5-6 $\upmu$m region, absent in most harmonic spectra but observed in AIBs (Jourdain de Muizon et al., 1986; Sloan et al., 1997). This feature may originate from the redshift of the C-H stretching band in superhydrogenated PAHs due to anharmonic effects (Mackie et al., 2018a; Yang et al., 2020; Sandford et al., 2013), with additional redshift and intensity enhancement likely driven by temperature effects.
@@
Figure 7: IR spectra of a mixture of 1 704 neutral PAHs computed using MLMD at 50 K (blue), 300 K (green), and 600 K (red). For comparison, the mixture of scaled harmonic spectra from PAHdb, are shown in gray. Each spectrum is normalized to its respective maximum intensity.
IV Conclusions

### v3

# Computing Anharmonic Infrared Spectra of Polycyclic Aromatic Hydrocarbons Using Machine Learning Molecular Dynamics

@@ Abstract
We introduce a machine learning molecular dynamics (MLMD) approach to calculate the anharmonic infrared (IR) absorption spectra of polycyclic aromatic hydrocarbons (PAHs), key carriers of interstellar aromatic IR bands. This method accounts for temperature effects in a molecule-specific way and achieves accuracy comparable to conventional quantum chemical calculations at a fraction of the cost, scaling linearly with system size. We applied MLMD to calculate the anharmonic spectra of 1,704 PAHs in the NASA Ames PAH IR Spectroscopic Database with up to 216 carbon atoms at different temperatures, demonstrating its capability for high-throughput spectral calculations of large molecular systems.
pubyear: 2025pagerange: Computing Anharmonic Infrared Spectra of Polycyclic Aromatic Hydrocarbons Using Machine Learning Molecular DynamicsReferences
1 Introduction
@@
In the mid-1970s, distinct infrared (IR) signals were identified in the universe, notably at wavelengths of 3.3, 6.2, 7.7, 8.6, 11.2, and 12.7 $\upmu$m. These signals were later linked to the vibrations of C-H and C-C bonds in aromatic molecules, known as aromatic IR bands (AIB) (Leger & Puget, 1984; Allamandola et al., 1985). Such emissions are primarily attributed to polycyclic aromatic hydrocarbons (PAHs) after the absorption of ultraviolet (UV) photons. PAHs are recognized as key contributors to the evolution of the interstellar medium, influencing processes such as gas heating, ionization balance, and star formation (Tielens, 2008). Determining the population, composition, size distribution, charge state, and chemical structures of interstellar PAHs provides valuable insight into the physical conditions of their host galaxies. However, this remains a significant challenge due to the structural diversity of PAHs. Variations in carbon-ring arrangements, side groups, substitutions, charge states, and isotopologues lead to distinct IR emission characteristics, each governed by specific vibration modes (Peeters et al., 2021).
@@
In order to tackle these challenges, quantum chemical calculations (QCCs) are now an indispensable aid in AIB analysis. Using QCCs, publicly available spectral databases have been established for AIB analysis, such as the NASA Ames PAH IR Spectroscopic Database (PAHdb) (Mattioda et al., 2020; Bauschlicher Jr. et al., 2018; Boersma et al., 2014). The availability of these databases enables extensive data-driven analysis of AIBs, providing a more robust interpretation of their features, see, e.g., (Boersma et al., 2013, 2015; Maragkoudakis et al., 2020; Sadjadi et al., 2015; Cami, 2011; Shannon & Boersma, 2019; Ricca et al., 2021). Furthermore, these databases have facilitated the analysis of astrochemical components using machine learning (ML) models, demonstrating significant computational efficiency (Kovács et al., 2020) and helping to uncover relationships between molecular structures and spectra (Meng et al., 2021, 2023). Despite these advancements, current theoretical spectral databases face three primary limitations:
@@
  • Harmonic oscillator approximations: Most spectra neglect anharmonic vibrations, leading to significant discrepancies with experimental observations (Mackie et al., 2015, 2018a).
  • Neglect of molecule-specific thermal effects: Most theoretical spectra in databases such as PAHdb are calculated at the ground state, without accounting for temperature effects. Consequently, emission features derived from these absorption spectra involve significant approximations.
  • Underrepresentation of large PAHs: Although PAHs with 50 or more C atoms are considered more likely AIB carriers (Sellgren, 1984; Allamandola et al., 1989), they are significantly underrepresented in current databases.
@@
These challenges arise mainly from the high computational cost of QCCs. In particular, the computation time for anharmonic molecular spectra using second-order vibrational perturbation theory (VPT2) is prohibitively long, typically only accommodating molecules with $N_{\text{C}}<25$, where $N_{\text{C}}$ represents the number of carbon atoms in the system (Mackie et al., 2016; Esposito et al., 2024a; Maltseva et al., 2016, 2018; Lemmens et al., 2019). The cost of incorporating temperature effects into these calculations using the Wang-Landau random walk algorithm is even more prohibitive, being orders of magnitude higher due to the need for extensive sampling (Chen, 2018; Chen et al., 2019). Facing the vast diversity of PAH species, these limitations significantly hinder the accurate interpretation of observational data, posing substantial challenges in the era of advanced observational facilities such as the James Webb Space Telescope (JWST) (Boersma et al., 2023; Ricca et al., 2024).
@@
As a result, there is an urgent need for a cost-effective and accurate method that incorporates both anharmonic and temperature effects in the construction of spectral databases to improve our understanding of the origin of AIBs. In this work, we propose a ML-based molecular dynamics (MLMD) approach for computing PAH IR absorption spectra. This method replaces computationally expensive electronic structure calculations with far more efficient ML-based ones. More importantly, it explicitly accounts for anharmonic vibrations and temperature effects. We show that the MLMD approach can reproduce experimental spectra with accuracy comparable to state-of-the-art VPT2 QCCs, while significantly reducing computational time, achieving a scaling approximately linear in $N_{\text{C}}$.
2 Methods
@@
Figure 1: Schematic representation of the workflow for computing anharmonic IR absorption spectra.
@@
2.1 Training Data
The chemical structures of 687 neutral PAH molecules were extracted from the theoretical dataset of PAHdb (v3.2, Mattioda et al. (2020)) to construct the training datasets. The selected PAHs encompass a diverse array of chemical structures that are of astronomical significance. Classical MD simulations were performed using a custom code to generate random configurations of the atoms in each molecule at 300 K, using the adaptive interatomic reactive empirical bond order potential (Stuart et al., 2000). In these simulations, the molecules reached thermal equilibrium in 200 ps using a canonical Nosé-Hoover thermostat, with a time step of 0.5 fs. Ten extended structures were extracted from the atomic trajectories of each molecule. As these geometries do not necessarily fall near equilibrium configurations, random Gaussian perturbations were applied to the ground-state atomic coordinates to create 15 more configurations for each molecule to improve the generalizability of the model. Finally, in total, 17,175 configurations (also called molecular conformations) were generated. Our training dataset includes 350 molecular conformations of superhydrogenated PAHs, out of a total of 17,175 conformations.
Subsequently, single-point electronic structure QCCs were performed for each of these 17,175 configurations to calculate the potential energy, the forces in the nuclei, and the dipole $\mathbf{p}$, within the framework of density functional theory (DFT) at the B3LYP/4-31G level, as implemented in the Gaussian 16 software package (Frisch et al., 2016). This level of theory offers a favorable balance between accuracy and computational efficiency for large PAHs (Stephens et al., 1994), given its extensive application in PAH IR spectrum studies (Bauschlicher Jr. & Bakes, 2000; Ricca et al., 2012). We note that the basis set does not include polarization functions, which may introduce more uncertainties than larger basis functions (e.g 6-31g*) for PAHs containing nitrogen (Ricca et al., 2021, 2024). However, this choice was made to ensure computational efficiency, which is crucial for large-scale sampling. Through these calculations, we generated two datasets for training the NNFF and EPNN models, respectively. The first dataset comprises 17 175 energy and 2 274 825 force data points, while the second dataset includes 17 175 molecular dipole moments.
2.2 NNFF
@@
The training data of NNFF were computed at the B3LYP/4-31G level of DFT, which is known to systematically overestimate the bond strength. To correct for this, we applied a scaling factor of 0.9578 to the MLMD-predicted vibrational frequencies, following the approach proposed by Bauschlicher & Langhoff (1997).
2.3 EPNN
@@
2.4 Anharmonic IR spectrum
To compute the anharmonic IR absorption spectrum of a PAH, MD simulations were performed to simulate its finite-temperature dynamics based on atomic forces predicted by the trained NNFF. We used the MD implementation in the Atomic Simulation Environment (ASE) (Larsen et al., 2017). Initially, the molecule reaches thermal equilibrium in 200 ps in a canonical ensemble (NVT) at a target temperature controlled by the Nosé-Hoover thermostat, with a timestep of 0.5 fs. The system was then further simulated in a microcanonical ensemble (NVE) for 200 ps, recording atomic configurations at every 1.0 fs during vibration.
These configurations serve as input to the EPNN, which is used to compute $\mathbf{p}(t)$ at each time step of the vibration. The dipole time autocorrelation function $\langle\mathbf{p}(0)\cdot\mathbf{p}(t)\rangle$, characterizing the time evolution of $\mathbf{p}$, is then subjected to a Fourier transform to obtain the absorption cross-section $\sigma$ (in electrostatic unit) (Berens & Wilson, 1981):
@@ 2.4 Anharmonic IR spectrum
where $\omega$ denotes the angular frequency, $k_{\text{B}}$ represents the Boltzmann constant, $\hbar$ is the reduced Planck constant, $c$ is the speed of light, $T$ indicates the temperature, $t$ is time, and $D(\omega)$ is the quantum correction factor, which is set to 1.0 in this work. Further explanation of this equation is provided in the Supplementary Information.
$\sigma$ is connected to the laboratory-measured IR intensity $I$ as:
[…]
where $N_{A}$ is Avogadro’s number. To compare $I$ with the MLMD computed $\sigma$, we applied Lorentzian broadening to the experimental spectrum with a full width at half maximum (FWHM) of $2\text{cm}^{-1}$. While the error values vary with the FWHM choice (see Supplementary Information), our key conclusion remains unaffected by this selection.
2.5 Reference methods for comparison
@@
The 49 PAHs for which the experimental spectra are available in PAHdb (v3.2), ranging from 10 to 50 carbon atoms, exhibit a variety of chemical structures, as detailed in the Supplementary Information. The experimental IR spectra were obtained using matrix isolation techniques at low temperatures (Hudgins & Allamandola, 1995; Hudgins & Sandford, 1998a, b, c; Hudgins & Allamandola, 1999; Mattioda et al., 2003, 2005, 2014, 2017). To facilitate comparison with experimental data, we employed theoretical spectra computed using MLMD at an effective temperature of 50 K, which accounts for molecular vibrations under experimental conditions (Esposito et al., 2024b). Stochastic uncertainty analysis confirms size-independent prediction stability across PAHs, and the model inherently supports non-planar geometries (see, Supplementary Information).
3 Results and discussion
@@
Figure 3: Comparison of IR spectra for five PAHs computed using the MLMD at 50 K(left, red), DFT harmonic (middle, green), and DFT anharmonic (right, blue) methods, alongside the experimentally measured spectra (black). The intensities are normalized to the maximum value in each spectrum.
@@
Figure 4 presents an overview of the error distribution in predicting the IR spectra of the 49 PAHs relative to the experimental data, for each of the three computational methods. The results clearly demonstrate that the harmonic spectra exhibit significantly poorer agreement with the experimental data compared to the anharmonic spectra, consistent with the findings of previous studies (Mackie et al., 2015; Maltseva et al., 2016; Lemmens et al., 2019). The precision of the MLMD and DFT anharmonic methods is comparable, with mean RMSE values of 0.0439 and 0.0434, respectively. For PAHs with fewer than 12 carbon atoms, the DFT anharmonic method generally outperforms MLMD. However, as the size of the PAHs increases, MLMD tends to provide comparable predictions than DFT anharmonic in terms of normalized spectra.
Figure 4: Distribution of RMSE values for spectra predicted by the three computational methods, compared to the experimental spectra of the 49 PAHs. $N_{\text{C}}$ represents the number of carbon atoms in each molecule. Note that the anharmonic DFT calculation is not performed for large PAHs (Section 2.5).
To further clarify the differences between MLMD and DFT in specific vibrational regions, a mode-by-mode error analysis is presented in Table 2. The results show that MLMD achieves higher accuracy in the 2.6–3.5, 5.2–5.8, and 6.0–10.0 µm regions, corresponding to C–H stretching, combination bands, and C–H in-plane (IP) bending modes, respectively. In contrast, for out-of-plane (OOP) bending modes in the 10.0–15.0 µm region, the DFT-based anharmonic method demonstrates better agreement with experimental data.
@@
Table 2: Mean RMSE values of IR spectra predicted by MLMD and DFT anharmonic methods, with respect to experimental spectra of the 26 PAHs across spectral regions of key vibrational modes.
Wavelength (µm) Vibration mode DFT anharm. MLMD
2.6-3.5 CH stretching 0.0534 0.0403
5.2-5.8 Combination 0.0103 0.0044
6-10 CH IP bending 0.0301 0.0273
10-15 CH OOP bending 0.0892 0.1067
@@
To demonstrate the precision, we compared theoretical anharmonic spectra with experimental IR spectra of two PAHs, pyrene and benzo[$k$]fluoranthene, measured at different temperatures. The higher temperature gas-phase experimental spectra have been taken from the NIST Chemistry WebBook database (Linstrom & Mallard, 2001), while the lower temperature spectra come from matrix isolation experiments of Hudgins & Sandford (1998a); Hudgins & Allamandola (1999). We note that the gas-phase IR spectra do not include explicitly stated temperature values, but prior studies (e.g. Mackie et al. (2015, 2016)) assumed an approximate measurement temperature of about 300 K. Figure 6 shows the NIST spectra (gray curves), referred to as $T_{\text{high}}$, and the low-temperature spectra (black curves), referred to as $T_{\text{low}}$. Comparison of the two sets of experimental spectra measured at different temperatures reveals that thermal effects lead to notable changes in spectral characteristics, such as band broadening and peak shifts (indicated by dashed lines), which aligns with previous observations of Chen (2018).
The two pairs of experimental spectra shown in the top panels of Figure 6 a and b are compared with three theoretical spectra in the bottom panels, calculated using MLMD at 300 (red) and 50 K (blue), as well as by the anharmonic DFT method (green). Compared with the 50 K MLMD results, the temperature-induced band broadening and peak shift behaviors are better captured by the MLMD results at 300 K, with relatively low RMSE values of 0.0565 and 0.0728 for the two molecules. In contrast, the agreement with the gas-phase experimental spectra is poorer for the 50 K results.
@@
Despite the growing need to incorporate anharmonicity and temperature effects into astronomical spectroscopic analyses, driven by high-resolution infrared observations in the JWST era (Peeters et al., 2021), only a limited number of theoretical anharmonic spectra for PAHs, likely fewer than 100, are available in the literature. Typically, these studies focus on just 2-5 PAH species, most of which contain fewer than 20 carbon atoms (Esposito et al., 2024c, d; Mackie et al., 2022, 2021). This narrow scope makes it difficult to draw general conclusions, especially given the vast diversity of PAH chemical structures. Such limitations hinder the accurate interpretation of AIBs and are mainly due to the high computational cost of traditional QCCs. To address this gap, we have computed the anharmonic spectra at 50, 300, and 600 K for 1,704 molecules from version 3.20 of PAHdb, which includes neutral PAHs as large as 216 carbon atoms. This dataset is open to the public as described in the Section of Data Availability.
@@
This dataset is expected to be a valuable resource for the spectral decomposition of AIBs and for data-driven studies investigating the precise structure-spectral relationships of PAHs. In addition, it will aid in the exploration of temperature effects. For instance, Figure 7 displays the anharmonic IR spectra of a simple mixture of the 1 704 neutral PAHs at 50 K (blue), 300 K (green), and 600 K (red). These spectra represent the sum of individual PAH spectra and are compared to the ground-state harmonic spectra from PAHdb (black, with scaling factors). Notably, the IR signals are significantly enhanced at high temperatures, exhibiting stronger peaks and an elevated plateau. A redshift in peak positions is also observed at high temperatures as a general trend. Furthermore, some emission bands that are weak at lower temperatures become significantly stronger at 300 and 600 K. For example, a distinct band emerges in the 3.5–6 $\upmu$m region, absent in most harmonic spectra but observed in AIBs (Jourdain de Muizon et al., 1986; Sloan et al., 1997). This feature may originate from the redshift of the C–H stretching band in superhydrogenated PAHs due to anharmonic effects (Mackie et al., 2018a; Yang et al., 2020; Sandford et al., 2013), with additional redshift and intensity enhancement likely driven by temperature effects.
@@
Figure 7: Absorption spectra of a mixture of 1,704 neutral PAHs computed using MLMD at 50 K (blue), 300 K (green), and 600 K (red). The three MLMD spectra were first normalized with respect to the total intensity of the 50 K spectrum, while the scaled harmonic spectrum from PAHdb (black) was normalized to its own total intensity.
4 Conclusions